OpenStudy (anonymous):

help pls. -> which is greater, an acceleration from 25km/h to 30km/h or an acceleration from 96km/hto 100km/h if both occur during the same time?

OpenStudy (anonymous):

v=vo+at 30=25+at 30-25=at 5=at 100=96+at 100-96=at 4=at

OpenStudy (anonymous):

Even without speed equations you can see that the 25-30 one is 1km/h greater.